OpenStudy (sissyedgar):

What is the exact area of a circle with a diameter of 18 feet? A. 36π square feet B. 81π square feet C. 324π square feet D. 1296π square feet @StudyGurl14

OpenStudy (acxbox22):

A=(pi)r^2 D=2r 18=2r radius=??

OpenStudy (sissyedgar):

I tink its A. 36π

OpenStudy (acxbox22):

why do you think it is A?

OpenStudy (sissyedgar):

cuz 18*2=36

OpenStudy (studygurl14):

They give you a diameter of 18. you need to find the radius. The diamter is 2 times the radius, so you need to divide 18 by 2 to find teh radius. then plug it into the formula @acxbox22 gave

OpenStudy (acxbox22):

first what is the radius if we have a diameter of 18?

OpenStudy (sissyedgar):

9?

OpenStudy (acxbox22):

yes now plug 9 into A=pi(9^2)

OpenStudy (sissyedgar):

WAIT! its 81! i kno! :P

OpenStudy (acxbox22):